OpenStudy (anonymous):

solve: y^-1 * dy+y e^cosx * sinx * dx=0

OpenStudy (kira_yamato):

\[y = -\frac{1}{e^{\cos x}}+ C\]

OpenStudy (anonymous):

big thanks (:

OpenStudy (anonymous):

dividing by dx on both sides of the equations y^-1*(dy/dx)+ye^cosx*sinx=0 => y^-1(dy/dx) = -y e^cosx*sinx => dy/dx= -y^2 * e^cosx*sinx